# SKIM: Any-bit Quantization Pushing The Limits of Post-Training Quantization

### v1

@@ 4 Methodology
The entire process of our SKIM algorithm is detailed in Figure 2. It begins by allocating different bits to each channel using a greedy algorithm, based on their quantization errors. Next, weighted K-means clustering is applied to each channel, generating centroids and labels with the allocated bits. Finally, we incorporate the scaling vector and train it through an iterative optimization strategy, keeping the labels fixed to enable gradient-based optimization. The dequantized weights can be recovered from the final labels and centroids, along with the scaling vector. Full algorithm of our SKIM method can be found in Appendix A. In the following subsections, we will explain our Mixed Precision and Scaling Vector techniques, as well as the principles behind our objective selection, in detail.
@@ 4.2 Mixed Precision
  Input: Error Matrix $E$, Minimum Available Bit $b_{\text{min}}$, Maximum Available Bit $b_{\text{max}}$, Row Number $n$, Bit $bit$
  Initialize $\mathbf{b}=\{b_{\text{min}}\}_{n}\in\mathcal{R}^{n}$
@@
     $i=\arg\max_{0\leq i<n}\{E_{i,\mathbf{b}_{i}}|\ \mathbf{b}_{i}<b_{\text{max}}\}$
     $\mathbf{b}_{i}=\mathbf{b}_{i}+1$
  until $\sum_{i=1}^{n}\mathbf{b}_{i}\geq n\cdot bit$
Additionally, our works includes efforts to enhance efficiency in mixed precision. Regarding computation for K-means clustering included in both error matrix pre-computing and the actual quantization, we develop an efficient framework utilizing parallel execution and shared memory. Furthermore, to accelerate the packing and unpacking phases with mixed precision, we gather all rows with the same bit and pack or unpack them together, leveraging the massive computational capabilities of modern GPUs. You can found detailed information about the time efficiency of our quantization method in Appendix D.2.
@@ 4.3 Scaling Vector
  Input: Weight $W$, Iteration $I$ = 1
@@
        $loss.backward()$
@@ Appendix A Full Algorithm
The full algorithm for our SKIM method is illustrated in Algorithm 3. Note that we provide the simplified version with number of iterations equal to one. Our method includes three main steps: 1. compute the bit allocation (line 2-3); 2. apply channel-wise K-Means Clustering (line 5-7); 3. train the scaling vector (line 8-14).
@@
2:  $E=pre\_record(W)\in\mathbb{R}^{n\times(b_{\text{max}}-b_{\text{min + 1}})}$ // pre-record quantization errors with Equation 5
3:  $b=alloc\_bit(E,\ b_{\text{min}},\ b_{\text{max}},\ n,\ bit)\in\mathbb{R}^{n}$ // compute bit allocation with Algorithm 1
4:  $L=\mathbf{0}_{n\times m}\in\mathcal{R}^{n\times m}$
5:  for $i=1$ to $n$ do
6:     $L_{i,:}\leftarrow Kmeans(W_{i,:},\ weights=G_{i,:},\ n\_centriods=2^{b_{i}}).labels$ // optimize on grouping towards Equation 4
7:  end for
8:  Initialize $\boldsymbol{\alpha}=\mathbf{1}_{1\times m}\in\mathcal{R}^{1\times m}$
9:  repeat
10:     $c=calc\_centroids(W*\boldsymbol{\alpha}^{-1},L)$
11:     $W^{q}=replace(c,L)*\boldsymbol{\alpha}$// replace labels with centroids and unscale to reconstruct the weight
12:     $loss=Err(W,W^{q})$// calculate quantization error with Equation 5
13:     $loss.backward()$// optimize on scaling
14:  until Converged
15:  return $L$, $c$, $\alpha$

### v2

@@ 4 Methodology
The entire process of our SKIM algorithm is detailed in Figure 2. It begins by allocating different bits to each channel using a greedy algorithm, based on their quantization errors. If channel $i$ is allocated $b_{i}$ bits, $2^{b_{i}}$ centroids will be generated for it. Next, weighted K-means clustering is applied to each channel to calculate the centroids and labels, using the allocated bits as a constraint. Finally, we incorporate the scaling vector and train it through an iterative optimization strategy. We keep the labels fixed to enable gradient-based optimization, which means the vector only adjusts centroids. The dequantized weights can be recovered from the final labels and centroids, along with the scaling vector. Full algorithm of our SKIM method is illustrated in Appendix A. In the following subsections, we will explain our mixed precision and scaling vector techniques, as well as the principles behind our objective selection, in detail.
@@ 4.2 Mixed Precision
0:  Error Matrix $E$, Minimum Available Bit $b_{\text{min}}$, Maximum Available Bit $b_{\text{max}}$, Row Number $n$, Bit $bit$
0:  Bit Allocation $b$
  Initialize $b=\{b_{\text{min}}\}_{n}\in\mathcal{R}^{n}$
@@
     $i=\arg\max_{0\leq i<n}\{E_{i,b_{i}+1}-E_{i,b_{i}}|\ b_{i}<b_{\text{max}}\}$
     $b_{i}=b_{i}+1$
  until $\sum_{i=1}^{n}b_{i}\geq n\cdot bit$
  return $b$
Additionally, our works includes efforts to enhance efficiency in mixed precision. Regarding computation for K-means clusterings, we develop an efficient framework utilizing parallel execution and shared memory. Furthermore, to accelerate the packing and unpacking phases with mixed precision, we gather all rows with the same bit and process them together, leveraging the massive computational capabilities of modern GPUs. You can found details about the time efficiency of our method in Appendix D.2.
@@ 4.3 Scaling Vector
0:  Weight $W$, Iteration $I$ = 1
0:  Labels $L$, Centroids $c$, Scaling Vector $\alpha$
@@
        $loss.backward()$// optimize on scaling
@@ Appendix A Full Algorithm
The full algorithm for our SKIM method is illustrated in Algorithm 3. Note that we provide the simplified version with number of iterations equal to one. Our method includes three main steps: 1. compute the bit allocation (line 2-11); 2. apply channel-wise K-Means Clustering (line 12-15); 3. train the scaling vector (line 16-22). Additionally, error pre-recording (line 2-10) only needs to be executed once for each model, and all KMeans functions are accelerated using multi-processing and shared memory.
@@
2:  $E=[\ 0\ ]_{n\times(b_{\text{max}}-b_{\text{min}}+1)}\in R^{n\times(b_{\text{max}}-b_{\text{min}}+1)}$ // Error Matrix
3:  for $i=1$ to $n$ do
4:     for $\hat{b}=b_{\text{min}}$ to $b_{\text{max}}$ do
5:        $result\leftarrow Kmeans(W_{i,:},\ weights=G_{i,:},\ n\_centriods=2^{\hat{b}})$ // apply Kmeans($\cdot$) with actual configuration
6:        $l,c\leftarrow result.labels,\ result.centroids$
7:        $w^{q}=replace(c,l)$ // replace labels with corresponding centroids
8:        $L_{i,\hat{b}}=Err(W_{i,:},w^{q})$ // pre-record quantization error with Equation 5
9:     end for
10:  end for
11:  $b=alloc\_bit(E,\ b_{\text{min}},\ b_{\text{max}},\ n,\ bit)\in\mathbb{R}^{n}$ // compute bit allocation with Algorithm 1
12:  $L=\mathbf{0}_{n\times m}\in\mathcal{R}^{n\times m}$
13:  for $i=1$ to $n$ do
14:     $L_{i,:}\leftarrow Kmeans(W_{i,:},\ weights=G_{i,:},\ n\_centriods=2^{b_{i}}).labels$ // optimize on grouping towards Equation 4
15:  end for
16:  Initialize $\boldsymbol{\alpha}=\mathbf{1}_{1\times m}\in\mathcal{R}^{1\times m}$
17:  repeat
18:     $c=calc\_centroids(W*\boldsymbol{\alpha}^{-1},L)$
19:     $W^{q}=replace(c,L)*\boldsymbol{\alpha}$// replace labels with centroids and unscale to reconstruct the weight
20:     $loss=Err(W,W^{q})$// calculate quantization error with Equation 5
21:     $loss.backward()$// optimize on scaling
22:  until Converged
23:  return $L$, $c$, $\alpha$